OpenStudy (jamanorthi):

i really need help a farmer sent his two child out to count the number of ducks and cows in the field.leo counted 50 heads.amayrani counted 154 legs. how many of each kind of animals was counted?

OpenStudy (austinl):

\(d+c=50\) \(2d+4c=154\) \(d=ducks\\ c=cows\) System of equations problem.

OpenStudy (anonymous):

let no of cows=x no of ducks=y each creature will have only 1 head total no of heads=x+y=50 cows have 4 legs and ducks have 2 each so 4x+2y=154 2x+y=77 subtract one equation from another you get no of cows then no of ducks=50-no of cows

OpenStudy (jamanorthi):

i like he way austinL did it

OpenStudy (jamanorthi):

then what you do

OpenStudy (jamanorthi):

i know how to solve those two eqution

OpenStudy (jamanorthi):

thanks for your time bye

OpenStudy (austinl):

Ok, we have the two equations. \(d+c=50\) \(2d+4c=154\) We can solve it a couple of ways. The easiest would probably be to solve for c in the top equation and plug it into the bottom equation. \(c=50-d\) \(2d+4(50-d)=154\) \(-2d+200=154\) \(-2d=-46\) \(d=23\) You can then plug that into the original equation. \(d+c=50\) \(23+c=50\) \(c=27\)